# Diffractive vector meson photo-production in Oxygen–Oxygen and Neon–Neon ultra-peripheral collisions at the LHC

### v1

@@ I Introduction
Diffractive vector-meson photo-production can be studied at the LHC in ultra-peripheral collisions (UPCs) where one of the incoming ions emits a quasi-real photon that fluctuates into a vector meson or a quark-antiquark colour dipole and interacts through the exchange of at least two gluons [23] with the other incoming ion; for a review see e.g. Refs. [24, 25, 26]. Within the Good-Walker formalism [27, 28, 29], two characteristics of the gluon field, its average and its fluctuations, can be studied via coherent, respectively incoherent, photo-production of vector mesons. In the former, the photon interacts with the full colour field of the nucleus, while in the latter the interaction occurs with either a single nucleon or a hotspot within it [30]. Diffractive vector meson production at the LHC opens the door to study the gluon field of nuclear targets at very large energies, which allows for the search for gluon saturation, a state of the colour field of hadrons predicted to occur at small Bjorken-$x$ [31, 32], where gluon splitting and annihilation reach a dynamical equilibrium. Although a large amount of theoretical and experimental effort has been dedicated to the search for saturation effects, up to now the existence of this phenomenon, and at which energy it sets in, has not been experimentally established [33].
@@ III.1 Photo-nuclear production off oxygen
Figure 4, shows the predicted energy dependence for the two vector mesons, the two processes, and the two models of the nuclear structure. In this case, the different nuclear models predict a different normalisation. What is interesting is that the prediction for the coherent case using the Woods-Saxon model is below than the prediction using the alpha-cluster model, while the opposite is true for the case of incoherent production, where the Woods-Saxon shape predicts a larger cross section than the alpha-cluster model. This observation applies to the photo-nuclear production of both vector mesons. It is worth noting that the alpha-cluster model predicts a cross section for $\rho^{0}$ incoherent photo-nuclear production that decreases with energy, as observed for proton targets [56, 38], while the Woods-Saxon approach yields a flat or even slowly increasing cross section.

### v2

# Diffractive vector meson photo-production in oxygen–oxygen and neon–neon ultraperipheral collisions at energies available at the CERN Large Hadron Collider

@@ I Introduction
Diffractive vector-meson photo-production can be studied at the LHC in ultra-peripheral collisions (UPCs) where one of the incoming ions emits a quasi-real photon that fluctuates into a vector meson or a quark-antiquark colour dipole and interacts through the exchange of at least two gluons [53] with the other incoming ion; for a review see e.g. Refs. [19, 35, 7]. Within the Good-Walker formalism [31, 45, 42], two characteristics of the gluon field, its average and its fluctuations, can be studied, respectively, via coherent and incoherent photo-production of vector mesons. In the former, the photon interacts with the full colour field of the nucleus, while in the latter the interaction occurs with either a single nucleon or a hotspot within it [43]. Diffractive vector meson production at the LHC opens the door to study the gluon field of nuclear targets at very large energies, which allows for the search for gluon saturation, a state of the colour field of hadrons predicted to occur at small Bjorken-$x$ [32, 47], where gluon splitting and annihilation reach a dynamical equilibrium. Although a large amount of theoretical and experimental effort has been dedicated to the search for saturation effects, up to now the existence of this phenomenon, and at which energy it sets in, has not been experimentally established [46].
@@ III.1 Photo-nuclear production off oxygen
Figure 4, shows the predicted energy dependence for the two vector mesons, the two processes, and the two models of the nuclear structure. In this case, the different nuclear models predict a different normalisation. What is interesting is that the prediction for the coherent case using the Woods-Saxon model is below than the prediction using the alpha-cluster model, while the opposite is true for the case of incoherent production, where the Woods-Saxon shape predicts a larger cross section than the alpha-cluster model. This observation applies to the photo-nuclear production of both vector mesons. The difference in the predictions for the coherent cross section comes naturally from different normalization of thickness function for both nuclear models (see also e.g. Ref. [40] for the case of the total cross section). The difference in the predictions for the incoherent cross sections can be traced back to the fact that in the Woods–Saxon model, nucleons and hotspots can occupy a wider range of positions throughout the nucleus, leading to a large number of distinct configurations. In contrast, in the tetrahedral model, nucleons are grouped into relatively fixed clusters, which limits the number of possible configurations. This explanation also applies for the predictions using the Woods–Saxon and PGCM models of neon presented below. The alpha-cluster model not only produces a smaller incoherent cross section, but also predicts that the incoherent photo-nuclear production of $\rho^{0}$ decreases with energy, as observed for proton targets [11, 18], while the Woods-Saxon approach yields a flat or even slowly increasing cross section.